Castle, Inc., has no debt outstanding and a total market value of $250,000. Earnings before interest and taxes, EBIT, are projected to be $40,000 if economic conditions are normal. If there is strong expansion in the economy, then EBIT will be 20 percent higher. If there is a recession, then EBIT will be 20 percent lower. The firm is considering a debt issue of $105,000 with an interest rate of 4 percent. The proceeds will be used to repurchase shares of stock. There are currently 10,000 shares outstanding. The firm has a tax rate 35 percent. Assume the stock price remains constant.

a-1.
Calculate earnings per share (EPS) under each of the three economic scenarios before any debt is issued. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EPS
Recession $
Normal $
Expansion $


a-2.
Calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answers as a percent rounded to the nearest whole number, e.g., 32.)

Percentage changes in EPS
Recession %
Expansion %


b-1.
Calculate earnings per share (EPS) under each of the three economic scenarios assuming the company goes through with recapitalization. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EPS
Recession $
Normal $
Expansion $


b-2.
Given the recapitalization, calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
  

Percentage changes in EPS
Recession %
Expansion %
Answer #1

a-1.EPS is calculated as follows:

Recession

Normal

Expansion

EBIT

32,000

40,000

48,000

Less: Interest @4%

0

0

0

EBT

32,000

40,000

48,000

Less: Tax@35%

11,200

14,000

16,800

EFE

20,800

26,000

31,200

Number of Shares

10,000

10,000

10,000

EPS

2.08

2.6

3.12

% change in EPS:

Recession: (2.08-2.6)/2.6 = -20%

Expansion: (3.12-2.6)/2.6 = 20%

Stock Price = 250,000/10,000 = $25 per share

b-1: EPS is calculated as follows:

Recession

Normal

Expansion

EBIT

32,000

40,000

48,000

Less: Interest @4% on 105,000

4,200

4,200

4,200

EBT

27,800

35,800

43,800

Less: Tax@35%

9,730

12,530

15,330

EFE

18,070

23,270

28,470

Number of Shares

5,800

5,800

5,800

EPS

3.12

4.01

4.91

Shares repurchased = 105,000/25 = 4,200

Shares outstanding = 10,000 – 4200 = 5,800

b-2

% change:

Recession: (3.12 – 4.01)/4.01 = -22.19%

Expansion: (4.91-4.01)/4.01 = 22.44%